OpenStudy (anonymous):

Sam purchased 625 square feet of carpet for his bedroom, which is shaped as a square. What are the dimensions of his bedroom? A)Each side of the bedroom has a length of 25 feet. B)Each side of the bedroom has a length of 156.25 feet. C)Each side of the bedroom has a length of 390,625 feet. D)Each side of the bedroom has a length of 312.5 feet.

OpenStudy (anonymous):

hellllpppppp me someone plzzzzz

OpenStudy (anonymous):

Well, a square's area is given by A=s^2, so S=Sqrt(A)